# Fast Algorithms for Fourier extension based on boundary interval data thanks: The research is partially supported by National Natural Science Foundation of China (No. 12171455, RSF-NSFC 23-41-00002)

### v4

@@ Abstract
In this paper, we first propose a new algorithm for the computation of Fourier extension based on data near the two endpoints of the interval, which can obtain an effective Fourier series approximation for non-periodic functions. The algorithm calculates the extended part using the boundary interval data and then combines it with the original data to form the data of the extended function within a period. By testing the key parameters involved, their influence on the algorithm was clarified and an optimization setting scheme for the parameters was proposed. Compared with FFT, the algorithm only needs to increase the computational complexity by a small amount. Then, an improved algorithm for the boundary oscillation function is proposed. By refining the boundary interval grid, the resolution constant of the boundary oscillation function was reduced to approximately $1/4$ of the original method.
keywords:
Fourier Series, Fourier Extensions, Fourier Continuation, Resolution Constant
{AMS}
42A10, 65T40, 65T50
@@ 1 Introduction
Fourier series approximation has been widely used in various fields for smooth periodic functions. In many physical, engineering and exploration geophysical problems, such as analyzing periodic signals in electrical engineering, studying periodic motion in mechanics such as the oscillation of a pendulum under certain ideal conditions, or decomposing seismic/electromagnetic signals into different frequency components using Fourier series approximation [17, 19, 20]. The Fourier series approximation offers spectral convergence since more terms are added to the Fourier series, the approximation converges to the actual function in an efficient way in the frequency domain. Moreover, it can be computed numerically via the FFT (Fast Fourier Transform), which is a computationally efficient algorithm. However, the situation changes completely once the periodicity is lost. A jump discontinuity is imposed in the approximation at the domain boundaries, owing to the well-known Gibbs phenomenon. For instance, when we attempt to approximate a function that is not periodic over a given interval using Fourier series, we will observe this Gibbs phenomenon near the boundaries of the interval. Several methods have been developed to ameliorate the ill effects of the Gibbs phenomenon. For example, methods based on Gegenbauer polynomials, which have certain orthogonality properties that can be exploited to obtain better approximate functions [9, 10, 18]. Methods utilizing Padé approximations, which are rational function approximations, can also be used [7, 8]. And Fourier extension (or continuation) methods, which extend the function in a certain way to make it more suitable for Fourier-based approximation, are also among the approaches developed to address this problem [4, 6, 11, 12, 13, 14, 15, 21].
The idea of the existing discrete Fourier extension can be described as follows: Let $f$ be a non-periodic, smooth and sufficiently differentiable function over $[-1,1]$, which is sampled in a predefined set of points $\{t_{\ell}\}\subset[-1,1]$, the discrete Fourier extension $C(f)$ of $f$ is defined by
@@
To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system has to be solved. In [4], Boyd introduced truncated singular value decomposition (TSVD) to stabilize the calculation process, with an accuracy almost up to machine precision. A convergence analysis by Adcock et al. states that for analytic functions the method is at least superalgebraically convergent [2]. Utilizing randomized algorithms, Lyon provided a fast algorithm with $T=2$ [12]. In [14], Matthysen et al. presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation. This idea was further developed in [15] to form a fast algorithm for function approximation in 2D domains.
@@
Because the data of the two functions near the endpoints are the same, the same connection data can be used in the intervals $[-T,1)$ and $(1,T]$. An illustration of this process is shown in Fig. 1. This raises the issue of whether the data inside the interval really help in the calculation of the extension. In fact, a very efficient method, known as the FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations [5]. Recently, this method has also been extended to the approximation of two-dimensional functions[6]. In [16], the authors proposed a method to calculate the Fourier extension using the derivative on the boundary combined with a two-point Taylor interpolation polynomial. The extension part calculated by these two methods are more stable, but neither of them is superalgebraically convergent because polynomials need to be introduced. Judging from the numerical results given in the literature, the number of nodes required by these two methods was significantly higher than that of the previous method when obtaining the same accuracy. In [5], the authors identified a critical limitation in directly applying the original extension algorithm with relatively few nodes in the boundary interval: the inherent difficulty in achieving satisfactory approximation accuracy. Therefore, they turned to precede the FC(SVD) calculation by projecting them into the bases of orthogonal polynomials. However, existing theoretical analyses reveal that the accuracy of the Fourier extension is co-determined by two interdependent parameters –extension length and sampling ratio. In previous studies, researchers tended to increase the sampling ratio to improve accuracy. This is because in the full data algorithm, increasing the extension length does not seem to bring any benefits. In this study, we will point out that by increasing the extension length, the number of nodes required in the boundary interval can be effectively reduced. Specifically, we demonstrate that through optimal parameter combinations, an accurate extension computation can be achieved with a few boundary nodes deployment. The number of nodes required in the boundary interval can be fixed, and is only related to the desired accuracy.
In this paper, we first propose a new fast algorithm for the calculation of the Fourier extension, which has the following traits:
  • By optimizing the parameter settings, only a few nodes in the boundary interval are required to calculate the extension part; Thus, the method can be computed with FFT speed. Unlike the FC-Gram algorithm, the method does not require the introduction of polynomials. When the behavior of the boundary is not significantly better than that of the interior, the proposed algorithm requires significantly fewer nodes than the FC-Gram algorithm because it maintains the superalgebraic convergence property of the original algorithm.
@@ 2 Outline of the boundary interval algorithm
Let the positive integer $m_{\Delta}$ be the number of nodes taken in the boundary interval, and $S_{l},S_{r}$ be the set of corresponding subscripts,
$S_{l}=\{-M,-M+1,\ldots,-M+m_{\Delta}-1\},\quad S_{r}=\{M-m_{\Delta}+1,M-m_{\Delta}+2,\ldots,M\}.$ (5)
Let $T_{\Delta}>1$ and
$L_{\Delta}=2\times\left\lceil T_{\Delta}\times(m_{\Delta}-1)\right\rceil,\quad h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta},$ (6)
[…]
$J_{\Delta,1}=\{1,2,\ldots,m_{\Delta}\},~{}J_{\Delta,2}=\{\frac{L_{\Delta}}{2}+1,\frac{L_{\Delta}}{2}+2,\ldots,\frac{L_{\Delta}}{2}+m_{\Delta}\},~{}J_{\Delta}=J_{\Delta,1}\cup J_{\Delta,2}.$ (7)
@@
Figure 2: Calculation process of a periodic extension of $f(x)=x^{2}$.
For $j\in J_{\Delta}$, take
Then we calculate the extension function $g_{c}$ defined on the interval $[0,2\pi]$ of $g$, and obtain its values at $x_{j},j=1,2,\ldots,L_{\Delta}$ according to the method described in Section LABEL:sec31. Finally, we obtain the function $f_{c}$ with a period of $2+\lambda$, where
$\lambda=\frac{\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)}{M},$ (9)
and the data of the extension function $f_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\frac{L_{\Delta}}{2}-m_{\Delta}$ in one period $[-1,1+\lambda)$ is
Fig. 2 shows the basic process of the boundary interval algorithm using the function $f(x)=x^{2}$ as an example. Here, to make the graph clearer, we use a relatively small number of nodes. Fig. 2(a) shows the given data (indicated by the black dotted line) and boundary data selected for the calculation extension (indicated by the red circle). Fig. 2(b) shows the $g_{c}(x_{j})$ calculated from the boundary data. The data represented by blue circles represent the subsequent extension data required. Fig. 2(c) shows the data of $f_{c}$ in one period formed by combining the original data with the calculated extension data. It can be observed from the calculation process that the calculation of the extension part depends only on the boundary interval data.
@@ 3 The calculation of $g_{c}$
[…]
@@
Figure 3: The singular values of $A$ for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$.
@@
The situation of singular values reflects the ill-posedness of (15), regularization must be introduced. Referring to the treatment methods in [2], the truncated SVD least-squares solution to (15) can be given as
@@ 4 Computational complexity and resolution constant
The computational complexity of the proposed algorithm comprises three parts:
  • Calculation of $c_{\tau}$. The computational effort in this step is mainly concentrated on the singular value decomposition of matrix $A$, which is related to $m_{\Delta}$ and $\gamma$, and can be estimated as
@@
    • If $\gamma\geq 1$, the computational complexity is
@@
Next, we analyze the factors that affect these parameters in detail. First, we provide the following concept for the resolution constant, which comes from the literature [3]:
Definition 4.1.
Let $\left\{F_{M}\right\}_{M\in\mathbb{N}}$ be a sequence of approximations such that $F_{M}(f)$ depends only on the values of $f$ on an equispaced grid of $12M+$ 1 points. Let
$\mathcal{R}(\omega,\delta)=\min\left\{M\in\mathbb{N}:\|\exp(\text{i}\pi\omega t)-F_{M}(\exp(\text{i}\pi\omega t))\|_{\infty}\leq\delta\right\},\quad\omega>0,~{}~{}0<\delta<1,$
@@
$R(\omega,\delta)\sim r\omega,\omega\rightarrow\infty,$ (25)
@@
  • First, we need to consider the approximation ability of $g_{c}$ to the boundary interval data of the function $f$. For the function $f(x)=\exp(\text{i}\pi\omega t)$ on $[-1,1]$, since the calculation of $g_{c}$ is performed on $[0,2\pi]$ after intercepting the boundary interval data, the function corresponding to $g$ in (8) is
@@
In fact, the increase of $\gamma$ plays a role in reducing the discrete step length on $[-T,T]$, which can also be achieved by increasing $T$, as mentioned in the previous literature on the estimation of the resolution constant of the full-data algorithm [3]:
@@ 5.1 Testing the parameter $T_{\Delta}$
(a) $\gamma=1$, $M=500$, $m_{\Delta}=100$.
(b) $\gamma=1$, $M=1000$, $\omega=20$.
(c) $\gamma=2$, $M=500$, $m_{\Delta}=150$.
(d) $\gamma=2$, $M=1000$, $\omega=20$.
(e) $\gamma=1/2$, $M=500$, $m_{\Delta}=50$.
(f) $\gamma=1/2$, $M=1000$, $\omega=20$.
@@
$\gamma$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
@@
It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma_{\Delta}$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma_{\Delta}$, $M$, and $\omega$. For several different values of $\gamma_{\Delta}$, we give the approximate values of $\hat{T}_{1}$ (we tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than $131e-$ 13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1. For $\hat{T}_{2}$, we can estimate:
@@
From the test results above, we can see that when $\gamma_{\Delta}$ is fixed, the selection of $T_{\Delta}$ is crucial to the approximation accuracy. The approximation can obtain satisfactory results only when $T_{\Delta}$ is within the interval $[\hat{T}_{1},\hat{T}_{2}]$. The left endpoint $\hat{T}_{1}$ of the interval is determined by parameter $\gamma_{\Delta}$, whereas the right endpoint $\hat{T}_{2}$ is affected by parameters $\gamma_{\Delta}$, $\omega$ and $M$.
5.2 Testing the parameter $m_{\Delta}$

### v5

@@ Abstract
This paper presents a novel boundary-optimized fast Fourier extension algorithm for efficient approximation of non-periodic functions. The proposed methodology constructs periodic extensions through strategic utilization of boundary interval data, which is subsequently combined with original function samples to form an extended periodic representation. We develop a parameter optimization framework that preserves superalgebraic convergence while requiring only a few boundary node deployment, resulting in computational complexity marginally exceeding that of standard FFT implementations. Furthermore, we present an improved version of the algorithm tailored for functions exhibiting boundary oscillations. This variant employs grid refinement near the boundaries, which reduces the resolution constant to approximately one-fourth of that in conventional approaches. Comprehensive numerical experiments confirm the efficiency and accuracy of the proposed methods and establish practical guidelines for parameter selection.
MSC:
42A10 41A17 65T40 42C15
@@ 1 Introduction
Fourier series approximation has been widely used in various fields for smooth periodic functions. In many physical, engineering and exploration geophysical problems, such as analyzing periodic signals in electrical engineering, studying periodic motion in mechanics such as the oscillation of a pendulum under certain ideal conditions, or decomposing seismic/electromagnetic signals into different frequency components using Fourier series approximation SteinFA2003 ; Tikhonov1995 ; WangIP2011 . The Fourier series approximation offers spectral convergence since more terms are added to the Fourier series, the approximation converges to the actual function in an efficient way in the frequency domain. Moreover, it can be computed numerically via the FFT (Fast Fourier Transform), which is a computationally efficient algorithm. However, the situation changes completely once the periodicity is lost. A jump discontinuity is imposed in the approximation at the domain boundaries, owing to the well-known Gibbs phenomenon. For instance, when we attempt to approximate a function that is not periodic over a given interval using Fourier series, we will observe this Gibbs phenomenon near the boundaries of the interval. Several methods have been developed to ameliorate the ill effects of the Gibbs phenomenon. For example, methods based on Gegenbauer polynomials, which have certain orthogonality properties that can be exploited to obtain better approximate functions 2006Robust ; 1992On ; 2006Optimal . Methods utilizing Padé approximations, which are rational function approximations, can also be used 2000A ; Geer1995 . And Fourier extension (or continuation) methods, which extend the function in a certain way to make it more suitable for Fourier-based approximation, are also among the approaches developed to address this problem Boyd2002A ; Bruno2022 ; Huybrechs2010 ; Lyon2011 ; 2012Sobolev ; Matthysen2016FAST ; Matthysen2018 ; Webb2020 .
The idea of the existing discrete Fourier extension can be described as follows: Let $f$ be a non-periodic, smooth and sufficiently differentiable function over [1,1][-1,1], which is sampled in a predefined set of points $\{t_{\ell}\}\subset[-1,1]$, the discrete Fourier extension $C(f)$ of $f$ is defined by
@@
To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system has to be solved. In Boyd2002A , Boyd introduced truncated singular value decomposition (TSVD) to stabilize the calculation process, with an accuracy almost up to machine precision. A convergence analysis by Adcock et al. states that for analytic functions the method is at least superalgebraically convergent adcock2014numerical . Utilizing randomized algorithms, Lyon provided a fast algorithm with $T=2$ Lyon2011 . In Matthysen2016FAST , Matthysen et al. presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation. This idea was further developed in Matthysen2018 to form a fast algorithm for function approximation in 2D domains.
@@
Because the data of the two functions near the endpoints are the same, the same connection data can be used in the intervals [T,1)[-T,1) and $(1,T]$. An illustration of this process is shown in Fig. 1. This raises the issue of whether the data inside the interval really help in the calculation of the extension. In fact, a very efficient method, known as the FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations bruno2010high . Since then, a lot of research has been done on further improvement and application of the FC-Gram algorithm 2016An ; 2020A ; Bruno2022 . In plonka2018numerical , the authors proposed a method to calculate the Fourier extension using the derivative on the boundary combined with a two-point Taylor interpolation polynomial. The extension part calculated by these two methods are more stable, but neither of them is superalgebraically convergent. When the boundary interval is more volatile than the interior, the number of nodes required to achieve near-machine precision by these two methods was higher than that of the previous method. In bruno2010high , the authors identified a critical limitation in directly applying the original extension algorithm with relatively few nodes in the boundary interval: the inherent difficulty in achieving satisfactory approximation accuracy. To address this, they introduced a pre-processing step involving Gram polynomial basis generation. Existing theoretical analyses reveal that Fourier extension accuracy is co-determined by two interdependent parameters - extension length and sampling ratio. While previous studies prioritized increasing sampling ratios for accuracy improvement (as extension length adjustments offered limited benefits in full-data algorithms), this work demonstrates that strategically increasing extension length can effectively reduce boundary node requirements. Specifically, we establish that through optimal parameter combinations, accurate extension computation can be achieved with a few boundary node deployment, where the required node count becomes fixed and solely dependent on target accuracy.
This paper presents a novel fast Fourier extension algorithm with the following key innovations:
  • Parameter optimization enables extension computation using minimal boundary nodes while maintaining FFT-compatible efficiency. Unlike FC-Gram approaches, our method eliminates polynomial basis requirements.
@@ 2 Outline of the boundary interval algorithm
Let $m_{\Delta}\in\mathbb{N}^{+}$ denote the number of boundary interval nodes, with corresponding index sets
$S_{l}=\{-M,-M+1,\ldots,-M+m_{\Delta}-1\},$ (5)
$S_{r}=\{M-m_{\Delta}+1,M-m_{\Delta}+2,\ldots,M\}.$
The calculation of the extension part is performed in the interval $0,2\pi]$. Define $T_{\Delta}>1$ as the scaling ratio between the extended interval $0,2\pi]$ and the data-containing interval, analogous to parameter $T$ in the original full-data algorithm. Further determine the number of nodes $L_{\Delta}$ used in the extension calculation and the corresponding discrete node $\{x_{j}\}$, specifically
$L_{\Delta}=2\times\left\lceil T_{\Delta}\times(m_{\Delta}-1)\right\rceil,$ (6)
$h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta},$
where $\lceil\cdot\rceil$ is the rounding symbol. Then let
$J_{\Delta,1}=\{1,2,\ldots,m_{\Delta}\},~J_{\Delta,2}=\{\frac{L_{\Delta}}{2}+1,\frac{L_{\Delta}}{2}+2,\ldots,\frac{L_{\Delta}}{2}+m_{\Delta}\},$ (7)
$J_{\Delta}=J_{\Delta,1}\cup J_{\Delta,2}.$
They indicate the locations of the boundary data during the extension calculation process. For $j\in J_{\Delta}$, take
Then we calculate the extension function $g_{c}$ defined on the interval $0,2\pi]$ of $g$, and obtain its values at $x_{j},j=1,2,\ldots,L_{\Delta}$ according to the method described in Section 3. Finally, we obtain the function $f_{c}$ with a period of $2+\lambda$, where
$\lambda=\frac{\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)}{M},$ (9)
and the data of the extension function $f_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\frac{L_{\Delta}}{2}-m_{\Delta}$ in one period $1,1+\lambda)$ is
@@
Figure 2: Calculation process of a periodic extension of $f(x)=\exp(x)$ ($T_{\Delta}=4,M=32$).
Fig. 2 shows the basic process of the boundary interval algorithm using the function $f(x)=\exp(x)$ as an example. Here, to make the graph clearer, we use a relatively small number of nodes. Fig. 2(a) shows the given data (indicated by the black dotted line) and boundary data selected for the calculation extension (indicated by the red circle). Fig. 2(b) shows the $g_{c}(x_{j})$ calculated from the boundary data. The data represented by blue circles represent the subsequent extension data required. Fig. 2(c) shows the data of $f_{c}$ in one period formed by combining the original data with the calculated extension data. It can be observed from the calculation process that the calculation of the extension part depends only on the boundary interval data.
@@ 3 The calculation of $g_{c}$
$A_{l,k}$ $=$ $\left\{\begin{aligned} &\phi_{k}(x_{l}),&1\leq l\leq m_{\Delta},\\ &\phi_{k}(x_{\frac{L_{\Delta}}{2}+l-m_{\Delta}}),&m_{\Delta}<l\leq 2m_{\Delta},\end{aligned}\right$. (14)
$b_{l}$ $=$ $\left\{\begin{aligned} &\frac{1}{m_{\Delta}}f(x_{l}),&1\leq l\leq m_{\Delta},\\ &\frac{1}{m_{\Delta}}f(x_{\frac{L_{\Delta}}{2}+l-m_{\Delta}}),&m_{\Delta}<l\leq 2m_{\Delta},\end{aligned}\right$.
@@
Figure 3: The singular values of AA for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$.
@@
The situation of singular values reflects the ill-posedness of (15), regularization must be introduced. Referring to the treatment methods in adcock2014numerical , the truncated SVD least-squares solution to (15) can be given as
@@ 4 Computational complexity and resolution constant
The computational complexity of the proposed algorithm decomposes into three principal components:
  • Calculation of $c_{\tau}$. The computational effort in this step is mainly concentrated on the singular value decomposition of matrix $A$, which is related to $m_{\Delta}$ and $\gamma_{\Delta}$, and can be estimated as
@@
    • If $\gamma_{\Delta}\geq 1$, the computational complexity is
@@
Next, we analyze the factors that affect these parameters in detail. First, we provide the following concept for the resolution constant, which comes from the literature adcock2014parameter :
Definition 1
Let $\left\{F_{M}\right\}_{M\in\mathbb{N}}$ be a sequence of approximations such that $F_{M}(f)$ depends only on the values of $f$ on an equispaced grid of $2M+1$ points. For $\omega>0,~~0<\delta<1,$ let
$\mathcal{R}(\omega,\delta)=\min\left\{M\in\mathbb{N}:\|\exp(\text{i}\pi\omega t)-F_{M}(\exp(\text{i}\pi\omega t))\|_{\infty}\leq\delta\right\},$
@@
$R(\omega,\delta)\sim r\omega,\quad\omega\rightarrow\infty,$ (25)
@@
  • First, we need to consider the approximation ability of $g_{c}$ to the boundary interval data of the function $f$. For the function $f(x)=\exp(\text{i}\pi\omega t)$ on [1,1][-1,1], since the calculation of $g_{c}$ is performed on $0,2\pi]$ after intercepting the boundary interval data, the function corresponding to $g$ in (8) is
@@
In fact, the increase of $\gamma$ plays a role in reducing the discrete step length on [T,T][-T,T], which can also be achieved by increasing $T$, as mentioned in the previous literature on the estimation of the resolution constant of the full-data algorithm adcock2014parameter :
@@ 5.1 Testing the parameter $T_{\Delta}$
(a) $\gamma_{\Delta}=1$, $M=500$, $m_{\Delta}=100$.
(b) $\gamma_{\Delta}=1$, $M=500$, $\omega=20$.
(c) $M=500$, $m_{\Delta}=100$, $\omega=20$.
(d) $\gamma_{\Delta}=2$, $m_{\Delta}=100$, $\omega=20$.
@@
$\gamma$ 88 44 22 11 $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
@@
It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma_{\Delta}$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma_{\Delta}$, $M$, and $\omega$. For several different values of $\gamma_{\Delta}$, we give the approximate values of $\hat{T}_{1}$ (we tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than 1e131e-13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1. For $\hat{T}_{2}$, we can estimate:
@@
From the test results above, we can see that when $\gamma_{\Delta}$ is fixed, the selection of $T_{\Delta}$ is crucial to the approximation accuracy. The approximation can obtain satisfactory results only when $T_{\Delta}$ is within the interval $\hat{T}_{1},\hat{T}_{2}]$. The left endpoint $\hat{T}_{1}$ of the interval is determined by parameter $\gamma_{\Delta}$, whereas the right endpoint $\hat{T}_{2}$ is affected by parameters $\gamma_{\Delta}$, $\omega$ and $M$.